# Efficient parallel scheduling for sparse triangular solvers

### v1

@@ 6.2.2. Erdős–Rényi
We generated thirty $N\times N$ matrices of this type with $N=$ 100,000 and $p=10^{-4},5\cdot 10^{-4},2\cdot 10^{-3}$, ten of each given probability. An overview over some statistics of the matrices may be found in Table LABEL:table:erdos-renyi-graph-table.
@@ 6.2.3. Narrow bandwidth
Unlike the Erdős–Rényi random matrices, we let the lower triangular matrix entry $(i,j)$, with $i>j$, being independently non-zero with probability $p\cdot\exp((1+j-i)/B)$, moving the non-zero entries closer to the diagonal. The entry values were chosen as in Section 6.2.2.
We generated thirty $N\times N$ matrices of this type with $N=$ 100,000 and $(p,B)=(0.14,10)$, $(0.05,20)$, $(0.03,42)$, ten for each pair $(p,B)$. An overview over some statistics of the matrices may be found in Table LABEL:table:random-bandwidth-graph-table.
@@ 6.3. CPU architectures
  • Intel Xeon Gold 6238T processor (x86), with 192 GB memory and theoretical peak memory throughput of 140.8 GB/s and 22 cores on a single socket;
  • AMD EPYC 7763 processor (x86), with 1024 GB memory and theoretical peak memory throughput of 204.8 GB/s and 64 cores on a single socket;
  • Huawei Kunpeng 920-4826 (Hi1620) processor (ARM), with 512 GB memory and theoretical peak memory throughput of 187.7 GB/s and 48 cores on a single socket.
@@ 7.1. Overall performance
We present speed-ups of the forward-/backward-substitution algorithm based on parallel schedules compared to serial execution. The schedules of our proposed algorithms are benchmarked against those produced by the baseline methods, SpMP [PSSD14] and HDagg [ZCL+22]. The results, aggregated over the instances from the respective data set using the geometric mean of all pairs of runs, are displayed in Table 7.1. All experiments were conducted on the Intel x86 machine utilizing 22 cores.
On our main data set, SuiteSparse, the schedules generated by our Funnel Locking algorithm achieved a geometric-mean speed-up of 1.45×1.45$\times$1.45 × compared to SpMP and 3.24×3.24$\times$3.24 × compared to HDagg. The data shows a similarly large improvement on the Erdős–Rényi data set. Here, the $p$-ivotal path algorithm achieved the best results, with a speed-up of 1.62×1.62$\times$1.62 × over SpMP and 1.87×1.87$\times$1.87 × over HDagg.
The narrow bandwidth graphs paint a slightly different picture. Here, the best schedule was produced by SpMP, with our algorithms falling behind. This is because these matrices allow for significantly less parallelizaton by design; such a setting is not a good fit for our schedulers, which aim to utilize most of the cores. In particular, Table 7.1 shows that when only fewer cores are available, our algorithms again outperform the baselines on this data set.
@@
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set & (Funnel) Locking (Funnel) $p$-ivotal path SpMP HDagg
SuiteSparse 10.709.35 10.449.27 7.39 3.30
Erdős–Rényi 14.0814.09 14.9215.00 9.23 7.96
Narrow bandw. 2.792.40 3.252.53 4.37 1.09
Table 7.1. Geometric mean of speed-ups over serial execution of $p$-ivotal path and Locking with/without Funnel coarsening, compared to the baselines SpMP and HDagg on the Intel x86 machine using 22 cores taken over the data sets described in Section 6.2.
@@
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set (Funnel) Locking (Funnel) $p$-ivotal path SpMP HDagg
Narrow bandw. 4.16 / 3.87 4.60 / 4.36 3.67 1.96
Table 7.2. Geometric mean of speed-ups over serial execution of $p$-ivotal path and Locking with/without Funnel coarsening, compared to the baselines SpMP and HDagg on the Intel x86 machine using 8 cores taken over the narrow bandwidth data set.
We also include a performance profile [DM02] based on the data generated from the SuiteSparse data set in Figure 7.1. The closer the line is to the top left corner, the better and more consistent the algorithm is across the data set. This shows that our algorithms are not only faster in execution time on average but are so throughout the diverse SuiteSparse data set.
Figure 7.1. Performance profiles of various algorithms on the SuiteSparse data set evaluated on the Intel x86 machine using 22 cores. The x-axis represents the threshold and the y-axis is the proportion of runs that are within the threshold times fastest SpTRSV run on the respective matrix.
7.2. Fewer synchronization barriers
The result in Table 7.1 show that our schedulers can significantly outperform the synchronous state-of-the-art HDagg. A further analysis shows that this is in part due to a substantial reduction in the number of synchronization barriers required during execution, whilst still maintaining a good work balance. In particular, Table 7.2 shows the number of synchronization barriers relative to the number of wavefronts in our algorithms and HDagg. The data indicates a large, up to 14×14$\times$14 ×, reduction of number of synchronization barriers compared to the number of wavefronts. This is a reduction of up to $\times1$ compared to HDagg, which explains the significant speed-ups achieved by our methods.
@@ 7.2. Fewer synchronization barriers
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set (Funnel) Locking (Funnel) $p$-ivotal path HDagg
SuiteSparse 14.00 / 9.63 13.26 / 9.76 1.24
Erdős–Rényi 2.75 / 2.74 3.29 / 3.26 1.25
Narrow bandw. 4.53 / 4.01 6.08 / 4.01 1.10
Table 7.3. Geometric mean of the reduction of the number of synchronization barriers relative to the number of wavefronts of the matrix.
7.3. Comparison of our algorithms
To compare our techniques and algorithms, we compute the geometric-mean speed-ups over Serial. We present average and interquartile ranges in Figure 7.2. The data shows that the Locking algorithms slightly outperform the $p$-ivotal path ones on the SuiteSparse data set, while the $p$-ivotal path algorithms have an edge on the Erdős–Rényi and the narrow bandwidth data sets. We observe that applying our coarsening algorithm, Funnel, greatly improves results on the SuiteSparse and the narrow bandwidth data sets. However, on the less structured matrices in the Erdős–Rényi data set, coarsening shows no performance gain.
(a) SuiteSparse
(b) Erdős–Rényi
(c) Narrow bandwidth
Figure 7.2. Geometric mean and interquartile ranges of speed-ups of (Funnel) $p$-ivotal path and (Funnel) Locking relative to Serial grouped according to data set. Experiments were conducted on the Intel x86 machine using 22 cores.
7.4. Impact of reordering
We separately analyze the impact of the reordering step on the performance. Table 7.4 compares the speed-ups achieved by our algorithms with and without the reordering component from Section 5. The numbers show that reordering is indeed a valuable ingredient of our schedulers. The data also confirms that even without the reordering, the algorithms still notably outperform HDagg, which is the current state-of-the-art synchronous baseline, cf. Table 7.1. Note that the narrow bandwidth random graphs are close to sequential by design, allowing limited parallelization, and thus any valid schedule here already exhibits good locality without reordering; as such, the reordering technique is not beneficial in this case. How to best reorder to improve locality for SpTRSV is research topic that deserves further attention.
@@ 7.4. Impact of reordering
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, rowhead = 1, rowfoot = 0, Data set Funnel Locking Locking Funnel $p$-ivotal p. $p$-ivotal p.
SuiteSparse 10.70 / 7.86 9.35 / 7.36 10.44 / 7.42 9.27 / 7.03
Erdős–Rényi 14.08 / 8.91 14.09 / 8.95 14.92 / 9.10 15.00 / 9.10
Narrow bandw. 2.79 / 3.19 2.40 / 2.64 3.25 / 3.20 2.53 / 2.69
Table 7.4. Geometric mean of speed-ups relative to Serial of our algorithms with/without permuting the matrix data according to the computed schedule. Experiments were conducted on the Intel x86 machine using 22 cores.
7.5. Performance across different architectures
We show the performance gains of our algorithms over the different processors and architectures in Table 7.5. The data confirms that our algorithms consistently outperform the baselines across all considered architectures. We note that the improvement relative to Serial can be in a significantly different range due to the properties of the distinct architectures. SpMP is omitted for the ARM architecture because its implementation is x86-specific.
{tblr}
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, rowhead = 1, rowfoot = 0, Machine (Funnel) Locking (Funnel) $p$-ivotal path SpMP HDagg
Intel x86 10.70 / 9.35 10.44 / 9.27 7.39 3.30
AMD x86 5.99 / 5.40 5.80 / 5.42 4.39 2.12
Huawei ARM 9.30 / 8.89 9.07 / 8.80 n/a 2.03
@@ 7.6. Scaling with the number of cores
Another natural question is how our algorithms scale with a growing number of cores. To examine this, we illustrate the speed-ups (over serial execution) for different numbers of cores in Table 7.6. We note that this experiment was conducted on the AMD x86 machine as it has $64$ available cores on a single socket.
@@
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, rowhead = 1, rowfoot = 0, Algorithm 2 4 8 16 24 32 40 48 56 64
Funnel Locking 1.86 2.92 3.40 5.24 6.19 6.72 6.92 7.02 6.87 6.65
Funnel $p$-ivotal path 1.85 2.89 3.31 5.11 6.04 6.45 6.62 6.76 6.63 6.52
Table 7.6. Geometric mean of speed-ups relative to Serial of Funnel $p$-ivotal path and Funnel Locking for different number of cores on the AMD x86 machine taken over the SuiteSparse data set.
As one sees, additional cores have diminished or negative returns at the higher end of number of cores. A reason for this is the average wavefront size which is a proxy for the amount of parallelism available. If we split the SuiteSparse data set into groups according to their average wavefront size, we see that these groups scale to different number of cores, see Figure 7.3. This shows that our algorithms do scale if the matrices allow for it and that the number of cores is an important parameter.
Figure 7.3. Geometric mean speed-ups of Funnel $p$-ivotal path for different number of cores on the AMD x86 machine taken over the SuiteSparse data set categorized by average wavefront size.
@@ 7.7. Amortization of scheduling time
In this final section, we consider the gain of the different scheduling algorithms when the scheduling time is taken into account. We measure the amortization threshold as the ratio of the time it took to compute the schedule to the difference in execution times between the serial execution and the parallel execution using the respective schedule [^3] If the parallel execution is slower than the serial one, then the amortization threshold is defined as +$\infty$+ ∞.. The same metric was considered by Zarebavani et. al. [ZCL+22, §V.B] and expresses how often the schedule needs to be reused in order to justify the time spent on computing it. Table 7.7 presents the amortization threshold for several scheduling algorithms with the 25th percentile, median, and 75th percentile values shown for each algorithm.
@@
colspec = Q[l,m] || Q[si=table-format=3.2, table-number-alignment=center,c] | Q[si=table-format=4.2, table-number-alignment=center,c] | Q[si=table-format=4.2, table-number-alignment=center,c] , row1 = font=, guard, rowhead = 1, rowfoot = 0, cell62 = font=, cell63 = font=, cell64 = font=, Algorithm 25th percentile Median 75th percentile
Funnel Locking 740.96 870.26 1158.46
Funnel $p$-ivotal path 731.91 867.67 1047.27
Locking 83.58 111.36 176.16
$p$-ivotal path 52.14 63.92 120.73
SpMP 4.91 6.44 8.93
HDagg 489.88 1306.06 2065.96
@@
A noticeable gap exists between the funneled and un-funneled versions of both the $p$-ivotal path and Locking algorithms. The funneled versions involve coarsening, a process that is computationally expensive and results in higher scheduling times. However, this added cost is justified when schedules are reused more than 1000 times in magnitude. In such scenarios, the coarsening pays off by reducing execution time, leading to a greater overall benefit in repeated parallel execution performance.
SpMP consistently exhibits low scheduling times across all percentiles, resulting in exceptional amortization threshold. This reflects the strong engineering behind SpMP, which minimizes scheduling overhead compared to the other algorithms. Our scheduling algorithms are research prototypes and single threaded, and as such, their run time could likely be reduced with further engineering. For example, the matrices may be split into blocks, c.f. Section 1.2, and our algorithms can be applied to each triangular block in parallel. The resulting block schedules are then easily combined into one schedule for the whole matrix. In any case, in scenarios where schedules are reused often enough, our proposed algorithms in their current implementation already demonstrate significant reduction in overall execution time.

### v2

@@ 6.2.4. Erdős–Rényi
We generated thirty $N\times N$ matrices of this type with $N=$ 100,000 and $p=10^{-4},5\cdot 10^{-4},2\cdot 10^{-3}$, ten of each given probability. An overview over some statistics of the matrices may be found in Table LABEL:table:erdos-renyi-graph-table of the supplement.
@@ 6.2.5. Narrow bandwidth
We also create a data set of random matrices which are much harder to parallelize by design, but have good locality. Unlike the Erdős–Rényi random matrices, we let the lower triangular matrix entry $(i,j)$, with $i>j$, be independently non-zero with probability $p\cdot\exp((1+j-i)/B)$, moving the non-zero entries closer to the diagonal. The entry values were chosen as in Section 6.2.4.
We generated thirty $N\times N$ matrices of this type with $N=$ 100,000 and $(p,B)=(0.14,10)$, $(0.05,20)$, $(0.03,42)$, ten for each pair $(p,B)$. An overview over some statistics of the matrices may be found in Table LABEL:table:random-bandwidth-graph-table of the supplement.
@@ 6.3. CPU architectures
  • Intel Xeon Gold 6238T processor (x86), with 192 GB memory and theoretical peak memory throughput of 140.8 GB/s and 22 cores on a single socket; kernel version 5.14.0; GCC version 11.5.0;
  • AMD EPYC 7763 processor (x86), with 1024 GB memory and theoretical peak memory throughput of 204.8 GB/s and 64 cores on a single socket; kernel version 5.15.0; GCC version 11.4.0;
  • Huawei Kunpeng 920-4826 (Hi1620) processor (ARM), with 512 GB memory and theoretical peak memory throughput of 187.7 GB/s and 48 cores on a single socket; kernel version 5.15.0; GCC version 11.4.0.
@@ 7.1. Overall performance
We present speed-ups of the forward-/backward-substitution algorithm based on parallel schedules compared to serial execution. The schedules of our proposed algorithm are benchmarked against those produced by the baseline methods, SpMP [PSSD14] and HDagg [ZCL+22]. The results, aggregated over the instances from the respective data set using the geometric mean of all pairs of runs, are displayed in Table 7.1. All experiments were conducted on the Intel x86 machine utilizing 22 cores.
On our main data set, SuiteSparse, the schedules generated by our GrowLocal algorithm achieves a geometric-mean speed-up of 1.42×1.42$\times$1.42 × compared to SpMP and 3.32×3.32$\times$3.32 × compared to HDagg. We also see similar results on the two variations of the SparseSuite data set: on METIS, GrowLocal obtains a 1.70×1.70$\times$1.70 × and 1.77×1.$7$ geometric-mean speed-up to SpMP and HDagg, respectively, and on iChol, it achieves a 1.80×1.80$\times$1.80 × and 2.20×2.20$\times$2.20 × speed-up to SpMP and HDagg, respectively. This shows that GrowLocal indeed significantly outperforms the baseline algorithms on these application-based data sets.
The differences are even larger on the Narrow Bandwidth matrices: here GrowLocal achieves a 2.50×2.50$\times$2.50 × and 10.12×10.12$\times$10.12 × factor improvement compared to SpMP and HDagg, respectively. This indicates that in the more challenging cases when our DAGs are particularly hard to parallelize, GrowLocal is even more superior to the baselines.
Finally, on the Erdős–Rényi data set, the improvement is much smaller; since these DAGs are easier to parallelize, the differences between the algorithms become less relevant.
We also note that Funnel coarsening does not seem to further improve GrowLocal; we elaborate on this later in Section 7.3.
@@
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[si=table-format=1.2, table-number-alignment=center,c,m] | Q[si=table-format=1.2, table-number-alignment=center,c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set & GrowLocal Funnel+GL SpMP HDagg
SuiteSparse 10.79 10.19 7.60 3.25
METIS 15.93 15.40 9.35 9.00
iChol 15.10 14.84 8.36 6.87
Erdős–Rényi 12.75 12.66 9.38 8.44
Narr. bandw. 9.04 8.26 3.56 0.88
Table 7.1. Geometric mean of speed-ups over serial execution of GrowLocal with/without Funnel coarsening, compared to the baselines SpMP and HDagg on the Intel x86 machine using 22 cores taken over the data sets from Section 6.2.
We also include a performance profile [DM02] based on the data generated from the SuiteSparse data set in Figure 7.1. The closer the line is to the top left corner, the better and more consistent the algorithm is across the data set. This shows that our algorithm is not only faster in execution time on average but it is so throughout the diverse SuiteSparse data set.
Figure 7.1. Performance profiles of our algorithms on the SuiteSparse data set evaluated on the Intel x86 machine using 22 cores. The x-axis represents a threshold and the y-axis is the proportion of runs that are within this threshold times the fastest SpTRSV run on the respective matrix.
7.2. Fewer synchronization barriers
The results in Table 7.1 show that our scheduler can significantly outperform the synchronous state-of-the-art HDagg. A further analysis shows that this is in part due to a substantial reduction in the number of synchronization barriers required during execution, whilst still maintaining a good work balance. In particular, Table 7.2 shows the number of synchronization barriers relative to the number of wavefronts in our algorithm and HDagg. The data indicates a large, up to 14.99×14.$9$, reduction of number of synchronization barriers compared to the number of wavefronts on the SuiteSparse data set. This is a reduction of up to 12.07×12.07$\times$12.07 × compared to HDagg, which explains the significant speed-ups achieved by our methods. In general, we can observe a similar effect in the remaining data sets, but the difference is much smaller for Erdős–Rényi, and much higher for narrow bandwidth matrices.
@@ 7.2. Fewer synchronization barriers
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set GrowLocal Funnel+GL HDagg
SuiteSparse 14.99 17.09 1.24
METIS 16.55 21.83 2.39
iChol 18.91 22.86 1.62
Erdős–Rényi 2.93 2.99 1.25
Narrow bandw. 51.12 42.00 1.10
Table 7.2. Geometric mean of the reduction of the number of synchronization barriers relative to the number of wavefronts of the matrix within each data set from Section 6.2.
7.3. Impact of Funnel coarsening
The results in Table 7.1 show that the DAG coarsening approach does not allow to further improve the schedules developed by GrowLocal on most of the data sets. This is an interesting contrast to the HDagg baseline, where coarsening is also a key ingredient of the scheduler. This suggests that GrowLocal is already rather strong at exploring the DAG structure and exploiting locality, and hence, the advantages of the coarsening step cannot compensate for the loss of structure in the graph.
However, besides this negative result, the Funnel coarsener also has several benefits that make it interesting in its own right. Firstly, it allows to run GrowLocal on a much smaller DAG, and as a results, the combined running time of Funnel+GrowLocal is often lower than GrowLocal alone; we will quantify this later in Section 7.7. As such, Funnel+GrowLocal can be a more desirable alternative when the scheduling time is also a critical factor. Secondly, Funnel coarsening allows one to reduce the number of synchronization barriers even further: while GrowLocal achieves a 12.07×12.07$\times$12.07 × geo-mean reduction compared to HDagg, Funnel+GrowLocal together achieves a 13.76×13.76$\times$13.76 × geo-mean reduction of synchronization barriers, which is of independent interest.
7.4. Impact of reordering
We separately analyze the impact of the reordering step on the performance. Table 7.4 compares the speed-ups achieved by our algorithm with and without the reordering component from Section 5. The numbers show that reordering is indeed a valuable ingredient of our scheduler. The data also confirms that even without the reordering, the algorithm still outperforms HDagg notably, which is the current state-of-the-art synchronous baseline, cf. Table 7.1.
@@ 7.4. Impact of reordering
colspec = Q[l,m] || Q[si=table-format=2.2, table-number-alignment=center,c,m] | Q[si=table-format=2.2, table-number-alignment=center,c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set Reordering No Reordering
SuiteSparse 10.79 8.62
METIS 15.93 15.21
iChol 15.10 15.02
Erdős–Rényi 12.75 7.87
Narrow bandw. 9.04 6.96
Table 7.3. Geometric mean of speed-ups relative to Serial of GrowLocal with/without permuting the matrix data according to the computed schedule. Experiments were conducted on the Intel x86 machine using 22 cores.
7.5. Performance across different architectures
We show the performance gains of our algorithm over the different processors and architectures in Table 7.5. The data confirms that our algorithm consistently outperforms the baselines across all considered architectures. We note that the improvement relative to Serial can be in a significantly different range due to the properties of the distinct architectures. SpMP is omitted for the ARM architecture because its implementation is x86-specific.
@@ 7.5. Performance across different architectures
colspec = Q[l,m] || Q[si=table-format=2.2, table-number-alignment=center,c,m] | Q[si=table-format=1.2, table-number-alignment=center,c,m] | Q[si=table-format=1.2, table-number-alignment=center,c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, cell43 = guard, Machine GrowLocal SpMP HDagg
Intel x86 10.79 7.60 3.25
AMD x86 5.20 3.65 1.98
Huawei ARM 9.27 n/a 2.16
@@ 7.6. Scaling with the number of cores
Another natural question is how our algorithm scales with a growing number of cores. To examine this, we illustrate the speed-ups (over serial execution) for different numbers of cores in Table 7.6. We note that this experiment was conducted on the AMD x86 machine as it has $64$ available cores on a single socket.
@@
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, rowhead = 1, rowfoot = 0, Algorithm 4 16 32 48 56 64
GrowLocal 2.63 4.15 5.34 5.70 5.76 5.85
As one sees, additional cores have diminished or negative returns at the higher end of number of cores. A reason for this is the average wavefront size which is a proxy for the amount of parallelism available. If we split the SuiteSparse data set into groups according to their average wavefront size, we see that these groups scale to different number of cores, see Figure 7.2. This shows that our algorithm does scale if the matrices allow for it.
Table 7.5. Geometric mean of speed-ups relative to Serial of GrowLocal for different number of cores on the AMD x86 machine taken over the SuiteSparse data set.
Figure 7.2. Geometric mean speed-ups of GrowLocal for different number of cores on the AMD x86 machine taken over the SuiteSparse data set categorized by average wavefront size.
@@ 7.7. Amortization of scheduling time
In this section, we consider the gain of the different scheduling algorithms when the scheduling time is taken into account. We measure the amortization threshold as the following ratio [^6] If the parallel execution is slower than the serial one, then the amortization threshold is defined as +$\infty$+ ∞.:
$\frac{\text{scheduling\_time}}{\text{serial\_execution\_time}-\text{parallel\_execution\_time}}\>.$ (7.1)
The same metric was considered by Zarebavani et. al. [ZCL+22, §V.B] and expresses how often the schedule needs to be reused in order to justify the time spent on computing it. Table 7.7 presents the amortization threshold for GrowLocal and the two baselines, with the 25th percentile, median, and 75th percentile values shown for each algorithm. The results indicate that the amortization threshold of GrowLocal (both with and without Funnel) is of a similar magnitude to that of SpMP. HDagg has a significantly higher amortization threshold on the SparseSuite data set. However, we remark that, e.g., on the METIS pre-processed data set we observed a comparable median amortization threshold of $44.21$ for HDagg.
@@
colspec = Q[l,m] || Q[si=table-format=3.2, table-number-alignment=center,c] | Q[si=table-format=3.2, table-number-alignment=center,c] | Q[si=table-format=4.2, table-number-alignment=center,c] , row1 = font=, guard, rowhead = 1, rowfoot = 0, cell62 = font=, cell63 = font=, cell64 = font=, Algorithm Q25 Median Q75
GrowLocal 23.78 26.12 30.28
Funnel+GL 17.78 21.74 27.78
SpMP 3.65 5.51 8.41
HDagg 311.23 961.39 1848.80
@@
7.8. Block parallel scheduling
In order to further reduce the amortization threshold of GrowLocal, we now consider the effect of subdividing the matrix into blocks and applying the GrowLocal scheduler on each block in parallel, cf. Section 3.1. In Table 7.8, we record the effect of running GrowLocal with multiple scheduling threads on scheduling time, floating point operations per second, number of supersteps, and amortization threshold. We see that using multiple scheduling threads can lead to super-linear speed-up of scheduling time. This is because there are now several (long) egdes that never have to be considered, cf. Figure 3.1. We also see that the SpTRSV solve time is more significantly affected when using a higher number of scheduling threads. We note that these effects are heavily dependent on the matrix. Some matrices, such as ‘af_shell7’ and ‘bmwcra_1’ have approximately a $30\%$ performance drop in the SpTRSV solve time already when using just two blocks, whereas ‘bundle_adj’ and ‘Hook_1498’ are hardly affected even when using $16$ blocks. Despite the performance drops, we find a near linear decrease in the amortization threshold.
Table 7.8 indicates, for instance, that using $6$ scheduling threads is a noteworthy compromise. This lowers the median amortization threshold to $4.54$, which is smaller than that of SpMP, cf. Table 7.7, whilst maintaining a 1.05×1.05$\times$1.05 × speed-up over SpMP. This makes GrowLocal superior on both metrics simultaneously.
{tblr}
colspec = Q[c,m] || Q[si=table-format=2.2, table-number-alignment=center,c,m] | Q[si=table-format=1.2, table-number-alignment=center,c,m] | Q[si=table-format=1.2, table-number-alignment=center,c,m] | Q[si=table-format=2.2, table-number-alignment=center,c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Threads Sched.
Time Flops/s Supersteps Amort.
Threshold
1 1.00 1.00 1.00 26.12
2 2.01 0.89 1.47 13.59
4 4.11 0.79 1.99 6.91
6 6.28 0.74 2.35 4.54
8 8.34 0.70 2.66 3.48
16 17.06 0.57 3.84 1.78
22 23.43 0.52 4.53 1.31
Table 7.7. Geometric means of relative speed-up of scheduling time, relative decrease in double precision floating point operations per second, and relative increase of number of supersteps of GrowLocal compared to using just a single scheduling thread, that is, a single scheduling block, together with the median amortization threshold of GrowLocal on the SuiteSparse data set. The data was collected on the Intel x86 machine using 22 cores.